persons basal metabolic rate (BMR) is the rate at which energy is expended while resting in a neutraly and kilocalories (or Calories) per hour. temperate environment. A ty pical BMR is 6.62 106 May. Convert this BMR to unts of watts (a) watts (b) klocalories per hour kcalh Suppose a 1.00 kg objects gravitation potential energy is increased at g-9.80 mys) a rate equal to this typical BMR Find the rate of change of the objects height in m/s. (The acceleration of gravity s rate nvs Need Help?s 100-g aluminum cup contains 950 g of water in thermal equilibrium with the cup at 75 C. The minute. At what rabe is energy being removed7 Express your answer in watts combination of cup and water is cooled uniformly so that the temperature decreases by 1.9 C per Need Help?s
Answer #1

Part(a). We know that -

1 j/day = 1.1574×10^-5 watt

Then 6.62×10^6 j/day = 11.574×6.62 =76.61988watt

Part(b). For kcal/hr

We know that -

1 watt= 0.86kcal/hr

Then , 76.61988 watt = 0.86×76.61988 = 65.8930968kcal/hr 狀 n) dt: 76 61 988 Nm/ree , dh AS oo.freA_{ 76-6 lage herlar 76-61 9 88 ↑kr9.gmle-. | 2